Write the net ionic equation to show the formation of a precipitate (insoluble salt) when the following solutions are mixed. Write noreaction if there is no precipitate.

AgF(aq) and KBr(aq)

Express your answer as a chemical equation. Identify all of the phases in your answer. Enter noreaction if no precipitate is formed.
1
Expert's answer
2015-11-03T12:33:47-0500
Solution
AgF(aq) + KBr(aq) = AgBr(s) + KF(aq)
Ag+ + Br- = AgBr(s)

Answer
Ag+ + Br- = AgBr(s)